# Ground states of strongly-correlated materials on quantum computers using ab initio downfolding

### v1

@@ III.1 Antiferromagnetism in $\text{Ca}_{2}\text{CuO}_{3}$
Figure 2: Band structure of $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The spin-state occupation of the ground state of downfolded $\text{Ca}_{2}\text{CuO}_{3}$ (panel b) indicates antiferromagnetic behavior, with the dependence of the antiferromagnetic order parameter $m$ on the lattice size given in panel c.
@@
We solve for the electronic ground state of $\text{Ca}_{2}\text{CuO}_{3}$, by performing VQE simulations of a one-dimensional, one-band Hubbard model using the above parameters, for a $10\times 1$ lattice at half-filling, i.e. we perform a $20$-qubit simulation of a Hamiltonian with $37$ terms ($20$ terms corresponding to the chemical potential of the single band are discarded since we can define $\mu=t_{i\mathbf{R}i\mathbf{R}}=0$). We use ten layers of the NP ansatz to represent our wavefunction, corresponding to $580$ variational parameters, which leads to a VQE energy within $11$ meV of the DMRG energy, and with a fidelity $\mathcal{F}=|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}=99.8\%$. Moreover, we see in Fig. 2b that the VQE ground state shows clear antiferromagnetic behavior, as is known to be the case for this material. We quantify the antiferromagnetic behavior by computing the staggered magnetization [59]
$m=\frac{1}{N_{x}N_{y}}\left|\sum_{x=1}^{N_{x}}\sum_{y=1}^{N_{y}}(-1)^{x+y}\langle\psi|\hat{S}_{z}(x,y)|\psi\rangle\right|,$ (7)
where |$\ket{\psi}$ the optimized VQE wavefunction. We visualize in Fig. 2c the value of $m$ for different lattices $N\times 1$, as a function of $N$, for $N=3-11$ (for $N=2$ no antiferromagnetic behavior is present). We see that $m$ assumes significant values, even if one were to extrapolate to $N\rightarrow\infty$, suggesting persistent antiferromagnetic behavior in this system in the thermodynamic limit.
@@ III.2 Excitonic ground state in $\text{WTe}_{2}$
Here we downfold the electronic structure on the subspace of four bands around the Fermi level (red in Fig. 3a). We were not able to perfectly reproduce the Kohn-Sham band structure with our computed Wannier-interpolated bands, with small deviations persisting towards the band edges. However, since this region is far from the Fermi level, it does not affect the formation of bound excitons. The crossing of the second and third bands of the subspace near the Fermi level is immediately suggestive of the possibility of exciton formation. To verify this, we obtain the ground state of the downfolded four-band extended Hubbard model within this subspace, the parameters for which are given in SI Section S3. Given the insulating character of this material, we solve our Hamiltonian at full-filling of the lower two states for a $2\times 2$ lattice, corresponding to a $2\times 2\times 4\times 2=32$-qubit simulation of a Hamiltonian with $288$ terms. We limit the bond dimension to $\chi=512$, and we utilize twenty layers of the EP ansatz, which leads to $1,304$ variational optimization parameters. The obtained ground state energy from VQE is within $0.12$ eV of the DMRG value, while the ground state VQE wavefunction has a fidelity of $90.5\%$. The difference $-\delta n_{el}$ in the total number of electrons within the bands of the strongly-correlated ground state we obtain with VQE (and with DMRG), from the number of electrons in the bands within the conventional band-insulating state (i.e., full-filling of the two lower-lying bands within the subspace, while the two higher-lying bands are empty), is visualized in Fig. 3b. The first observation to make here is that the DMRG and the VQE predictions are in excellent agreement. Moreover, it becomes evident that in our reduced system of $16$ electrons ($4$ lattice sites, times $2$ spin directions, times $2$ fully occupied electronic bands when we initialize), a significant lack of electrons occurs in the two lower-lying bands, i.e., positively charged holes form in the ground state. At the same time, a very substantial electronic population resides in the bands above the Fermi level (i.e., in the “conduction” bands). This presence of electrons in the “conduction” bands and holes in the “valence” bands is a hallmark of excitonic insulating behavior. To further quantify the formation of exciton pairs, we compute the excitonic insulator order parameter [63]
@@
where $v,c$ denote valence and conduction states respectively, and $U^{\prime}$ the inter-orbital, on-site Coulomb repulsion between electrons. Here we define $U^{\prime}$ as the average on-site, inter-orbital repulsive interaction between electrons, as these terms are given in SI Section S3. For a $2\times 2$ lattice we find $\Delta=0.693$, in reasonable agreement to the order parameter obtained using DMRG $\Delta_{DMRG}=0.876$ on the same lattice. The fact that VQE perfectly reproduces the DMRG band occupations, however leads to a finite error in the calculation of the order parameter is in agreement with previous observations on the Hubbard model, where it was found that even high-fidelity solutions can struggle to capture correlation functions [64], although our additional step of an overlap-based optimization substantially improves their accuracy [32]. Overall, for a small-sized lattice representation of $\text{WTe}_{2}$, our combined ab initio downfolding/VQE approach captures the previously reported excitonic ground state of this system, entirely from first principles.
@@ III.3 Charge-ordered state in $\text{SrVO}_{3}$
The band structure of $\text{SrVO}_{3}$ is visualized in Fig. 4a. The Fermi level crosses the subspace of the three vanadium $d$-bands, and hence semi-local DFT suggests that this system is a regular metal. However, $\text{SrVO}_{3}$ is known to be a correlated metal and to exhibit charge order [37, 38]. In order to obtain these signatures of strong correlations with our VQE approach, we downfold the electronic structure onto the active space of the three bands crossing the Fermi level (visualized in red in Fig. 4a), and we obtain a three-band extended Hubbard model, the parameters for which are given in SI Section S3, and which are in close agreement to previous downfolding calculations for this system [48].
We perform VQE simulations for a $3\times 3$ lattice of this three-band Hubbard Hamiltonian with $423$ terms, corresponding to a simulation with $3\times 3\times 3\times 2=54$ qubits. We apply ten layers of the EP ansatz, corresponding to $1,168$ variational parameters, with a bond dimension of $\chi=512$. While we were unable to obtain a VQE ground state of this complex system with a fidelity higher than $76.8\%$, our solution has an energy which is only $0.2$ eV above that of the DMRG solution, and as we will see, reproduces key manifestations of strong electronic interactions. In order to quantify the charge ordering in the ground state wavefunction, we compute the charge disproportionation parameter
@@
where $\sum_{A},\sum_{B}$ the total charge in two sublattices, which we define as $A$ and $B$ respectively. The total charge within a sublattice is obtained as $\sum_{A,B}=\sum_{b}\sum_{x}\sum_{y}\sum_{\sigma}n(x,y,b,\sigma)$, with $b$ a band index and a site belonging to sublattice $A$ if $(x+y)\mod 2=0$, and to sublattice $B$ otherwise. We find significant charge disproportionation with an order parameter $\Phi=0.30$, in reasonable agreement with DMRG simulations on the same lattice ($\Phi_{DMRG}=0.41$), and on a larger $4\times 4$ lattice ($\Phi_{DMRG}=0.40$). In Fig. 4b we visualize the charge distribution of the ground state wavefunction obtained with DMRG on a $4\times 4$ lattice, within the first two bands of the subspace, as this case most clearly illustrates the significant charge ordering which occurs. However, even for the $3\times 3$ lattices studied within VQE, we make qualitatively very similar observations.
@@ IV Conclusions and outlook
In this work we have utilized a combined ab initio downfolding-VQE approach to describe the ground states of strongly-correlated materials. We apply our approach to the antiferromagnetic quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic insulator $\text{WTe}_{2}$, and the correlated metal $\text{SrVO}_{3}$, and we demonstrate that we correctly predict the key physics of these systems, while semi-quantitatively reproducing the ground state energy as obtained using DMRG. This demonstrates the potential of our approach to utilize current and emerging quantum computing technologies in order to accurately predict the ground-state properties of diverse strongly-correlated materials, entirely from first quantum mechanical principles.
Our work raises several questions and highlights the importance of pursuing various avenues of research. The variational approaches we employ here can struggle to quantitatively reproduce the true ground state energy in the case of systems downfolded on two-dimensional lattices, it will therefore be important to explore flexible ansätze as ADAPT-VQE [65] for such problems moving forward. This will be particularly important as we are moving towards the fault-tolerant era [33] and quantum computer architectures with more capabilities, making it possible to describe increasingly complex materials, using our and similar approaches. Moreover, while here we have employed a static approximation for the Coulomb interactions within the active space, it will be interesting to perform a rigorous treatment of the frequency-dependence, as has been described in recent works [49, 50]. Furthermore, here we have employed the DFT generalized gradient approximation as a starting point for downfolding; it will be interesting to explore the influence of the starting point moving forward, as it has been shown previously that optimal DFT starting points can result in improved descriptions of correlated states with higher levels of electronic structure calculations [66, 67]. Finally, while here we have accounted only for electronic screening on Coulomb interactions, lattice motions have recently been shown to be capable of substantially modifying the screening, particularly in polar materials [68, 69], which could lead to modified downfolded Hamiltonians and phase diagrams [70, 71]; we reserve a detailed discussion of these effects for a future study.

### v3

@@ III.1 Antiferromagnetism in $\text{Ca}_{2}\text{CuO}_{3}$
Figure 3: Band structure of $\text{Ca}_{2}\text{CuO}_{3}$ (panel a), with the DFT bands given in black and the Wannier-interpolated bands in red. The spin correlation function of the ground state of downfolded $\text{Ca}_{2}\text{CuO}_{3}$ (panel b) indicates clear antiferromagnetic behavior.
@@
We solve for the electronic ground state of $\text{Ca}_{2}\text{CuO}_{3}$, by performing VQE simulations of a one-dimensional, one-band Hubbard model using the above parameters, for a $10\times 1$ lattice at half-filling, i.e. we perform a $20$-qubit simulation of a Hamiltonian with $37$ terms ($20$ terms corresponding to the on-site potential of the single band are discarded since we can define $t_{i\mathbf{R}i\mathbf{R}}=0$), compared to $2.83\times 10^{10}$ terms in the full many-body Hamiltonian, and $4.02\times 10^{4}$ terms in the many-body Hamiltonian, when restricted within the active space. We use ten layers of the NP ansatz to represent our wavefunction, corresponding to $580$ variational parameters, corresponding to $290$ two-qubit gates. If we assume a $99.9$% fidelity for two-qubit gates on near-term hardware, this suggests a $74.8$% circuit fidelity for a single VQE energy measurement. The 1-norm of the downfolded Hamiltonian in this system is found to be $||H||_{1}=2.67\times 10^{2}$, which is much lower than typical values of full many-body Hamiltonians [69], thus indicating the scalability of our approach to fault-tolerant architectures. These parameters, which relate to the near-term and fault-tolerant simulation of $\text{Ca}_{2}\text{CuO}_{3}$, and also of the other materials studied here, are summarized in Table 2.
Our tensor network simulation of VQE leads to a ground state energy within $23$ meV of the DMRG energy, and with a fidelity $\mathcal{F}=|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2}=99.3\%$. Moreover, we compute the spin correlation function
$\langle C_{ij}\rangle=\langle S_{i}^{z}S_{j}^{z}\rangle-\langle S_{i}^{z}\rangle\langle S_{j}^{z}\rangle,$ (8)
and we plot $\langle C_{1j}\rangle$ in Fig. 3b. The alternating sign of the spin correlation function indicates clear antiferromagnetic behavior for $\text{Ca}_{2}\text{CuO}_{3}$, as expected, and we see that VQE is in near-perfect agreement to DMRG.
@@ III.2 Excitonic ground state in $\text{WTe}_{2}$
Here we downfold the electronic structure on the subspace of four bands around the Fermi level (red in Fig. 4a). We were not able to perfectly reproduce the Kohn-Sham band structure with our computed Wannier-interpolated bands, with small deviations persisting towards the band edges. However, since this region is far from the Fermi level, it does not affect the formation of bound excitons. The crossing of the second and third bands of the subspace near the Fermi level is immediately suggestive of the possibility of exciton formation. To verify this, we obtain the ground state of the downfolded four-band extended Hubbard model within this subspace, the parameters for which are given in Section C of the Appendix. Given the insulating character of this material, we solve our Hamiltonian at full-filling of the lower two states for a $2\times 2$ lattice, corresponding to a $2\times 2\times 4\times 2=32$-qubit simulation of a Hamiltonian with $288$ terms, compared to $6.29\times 10^{8}$ terms in the full many-body Hamiltonian, and $2.63\times 10^{5}$ terms in the many-body Hamiltonian when restricted within the active space. We limit the bond dimension to $\chi=512$, and we utilize twenty layers of the EP ansatz, which leads to $1,304$ variational optimization parameters, corresponding to $652$ two-qubit gates and a $52.1$% circuit fidelity, if one were to assume a $99.9$% fidelity for the individual gates. The 1-norm of the downfolded Hamiltonian is $||H||_{1}=3.31\times 10^{2}$, which here too is significantly smaller than values of the order to $10^{4}$, which are obtained even for small chemical systems [69]. The obtained ground state energy from VQE is within $68$ meV of the DMRG value, while the ground state VQE wavefunction has a fidelity of $96.2\%$. The difference $-\delta n_{el}$ in the total number of electrons within the bands of the strongly-correlated ground state we obtain with VQE (and with DMRG), from the number of electrons in the bands within the conventional band-insulating state (i.e., full-filling of the two lower-lying bands within the subspace, while the two higher-lying bands are empty), is visualized in Fig. 4b. The first observation to make here is that the DMRG and the VQE predictions are in excellent agreement. Moreover, it becomes evident that in our reduced system of $16$ electrons ($4$ lattice sites, times $2$ spin directions, times $2$ fully occupied electronic bands when we initialize), a significant lack of electrons occurs in the two lower-lying bands, i.e., positively charged holes form in the ground state. At the same time, a very substantial electronic population resides in the bands above the Fermi level (i.e., in the “conduction” bands). This presence of electrons in the “conduction” bands and holes in the “valence” bands is a hallmark of excitonic insulating behavior. To further quantify the formation of exciton pairs, we compute the excitonic insulator order parameter [73]
@@
where $v,c$ denote valence and conduction states respectively, and $U^{\prime}$ the inter-orbital, on-site Coulomb repulsion between electrons. Here we define $U^{\prime}$ as the average on-site, inter-orbital repulsive interaction between electrons, as these terms are given in Section C of the Appendix. For a $2\times 2$ lattice we find $\Delta=0.379$, which is underestimated compared to the order parameter obtained using DMRG $\Delta_{DMRG}=0.640$ on the same lattice. The fact that VQE perfectly reproduces the DMRG band occupations, however leads to a finite error in the calculation of the order parameter is in agreement with previous observations on the Hubbard model, where it was found that even high-fidelity solutions can struggle to capture correlation functions [74], although our additional step of an overlap-based optimization substantially improves their accuracy [37]. Overall, for a small-sized lattice representation of $\text{WTe}_{2}$, our combined ab initio downfolding/VQE approach captures the previously reported excitonic ground state of this system, entirely from first principles.
@@ III.3 Charge-ordered state in $\text{SrVO}_{3}$
The band structure of $\text{SrVO}_{3}$ is visualized in Fig. 5a. The Fermi level crosses the subspace of the three vanadium $d$-bands, and hence semi-local DFT suggests that this system is a regular metal. However, $\text{SrVO}_{3}$ is known to be a correlated metal and to exhibit charge order [42, 43]. In order to obtain these signatures of strong correlations with our VQE approach, we downfold the electronic structure onto the active space of the three bands crossing the Fermi level (visualized in red in Fig. 5a), and we obtain a three-band extended Hubbard model, the parameters for which are given in Section C of the Appendix, and which are in close agreement to previous downfolding calculations for this system [50].
We perform VQE simulations for a $3\times 3$ lattice of this three-band Hubbard Hamiltonian with $423$ terms, compared to $7.34\times 10^{9}$ terms in the full many-body Hamiltonian, and $2.13\times 10^{6}$ terms in the many-body Hamiltonian, when restricted to the active space. In the compressed representation, our system is described using $3\times 3\times 3\times 2=54$ qubits. We apply ten layers of the EP ansatz, which requires to $1,168$ variational parameters, with a bond dimension of $\chi=512$. This corresponds to $584$ two-qubit gates, and assuming a fidelity of $99.9$% for each one, suggests a $55.8$% circuit fidelity on near-term hardware. The 1-norm of the downfolded Hamiltonian is $||H||_{1}=2.315\times 10^{3}$, which makes this system somewhat more challenging compared to the ones compared previously in terms of its potential simulation on fault-tolerant hardware. While we were unable to obtain a VQE ground state of this complex system with a fidelity higher than $31.8\%$, our solution has an energy which is only $18$ meV above that of the DMRG solution, and as we will see, qualitatively reproduces key manifestations of strong electronic interactions. In order to quantify the charge ordering in the ground state wavefunction, we compute the charge disproportionation parameter
@@
where $\sum_{A},\sum_{B}$ the total charge in two sublattices, which we define as $A$ and $B$ respectively. The total charge within a sublattice is obtained as $\sum_{A,B}=\sum_{b}\sum_{x}\sum_{y}\sum_{\sigma}n(x,y,b,\sigma)$, with $b$ a band index and a site belonging to sublattice $A$ if $(x+y)\mod 2=0$, and to sublattice $B$ otherwise. We find significant charge disproportionation with an order parameter $\Phi=0.21$, which is however overestimated compared to DMRG simulations on the same lattice ($\Phi_{DMRG}=0.12$). In Fig. 5b we visualize the charge distribution of the ground state wavefunction obtained with DMRG on a $4\times 4$ lattice, within the first two bands of the subspace, as this case most clearly illustrates the significant charge ordering which occurs. However, even for the $3\times 3$ lattices studied within VQE, we make qualitatively very similar observations.
@@ IV Conclusions and outlook
In this work we have utilized ab initio downfolding to generate compressed representations of strongly-correlated materials. We have shown that near-term quantum algorithms, such as the VQE, can produce the ground states of these compressed Hamiltonians and yields the expected behavior, at a modest computational cost and also in quantitative agreement with DMRG. We apply our approach to the antiferromagnetic quasi-1D cuprate $\text{Ca}_{2}\text{CuO}_{3}$, the excitonic insulator $\text{WTe}_{2}$, and the correlated metal $\text{SrVO}_{3}$, and we correctly predict the key physics of these systems, while semi-quantitatively reproducing the ground state energy as obtained using DMRG. This highlights the potential of our approach to utilize current and emerging quantum computing technologies in order to accurately predict the ground-state properties of diverse strongly-correlated materials, entirely from first quantum mechanical principles. Additionally, our classical simulation of VQE may be used as a pre-optimization step to the simulation of strongly-correlated materials on quantum hardware, hence providing an excellent starting point and minimizing the quantum resource cost [39, 75].
Our work raises several questions and highlights the importance of pursuing various avenues of research. The variational approaches we employ here can struggle to quantitatively reproduce the true ground state energy in the case of systems downfolded on two-dimensional lattices, it will therefore be important to explore flexible ansätze as ADAPT-VQE [76, 77] for such problems moving forward. This will be particularly important as we are moving towards the fault-tolerant era [38] and quantum computer architectures with more capabilities, making it possible to describe increasingly complex materials, using our and similar approaches. Moreover, while here we have employed a static approximation for the Coulomb interactions within the active space, it will be interesting to perform a rigorous treatment of the frequency-dependence, as has been described in recent works [51, 55]. Furthermore, here we have employed the DFT generalized gradient approximation as a starting point for downfolding; it will be interesting to explore the influence of the starting point moving forward, as it has been shown previously that optimal DFT starting points can result in improved descriptions of correlated states with higher levels of electronic structure calculations [78, 79]. Finally, while here we have accounted only for electronic screening on Coulomb interactions, lattice motions have recently been shown to be capable of substantially modifying the screening, particularly in polar materials [80, 81], which could lead to modified downfolded Hamiltonians and phase diagrams [82, 83, 84, 85]; we reserve a detailed discussion of these effects for a future study.
@@ V Acknowledgments
Appendix A Details of VQE optimization
Following Ref. [39], we start by optimizing for the ground state of the non-interacting system ($U=0$). We use this state |$\ket{\psi_{o}}$ to generate the parameterized quantum circuit
|$\ket{\psi_{\mathrm{PQC}}(\bm{\theta})}=U_{n}(\bm{\theta}_{n})...U_{1}(\bm{\theta}_{1})\ket{\psi_{o}}$ (11)
as the starting point for the optimization of the interacting case, representing the real material of interest. The initial parameters $\bm{\theta}$ are obtained randomly from a Gaussian distribution $\mathcal{N}(0,10^{-5})$ with zero mean and variance $\sigma^{2}=10^{-5}$. We perform an energy minimization, until one of three conditions is met: the energy tolerance (defined by the absolute difference between the energy at the final step and the penultimate step) reaches $10^{-7}$, the energy gradient reaches $10^{-6}$, or the optimization reaches $500$ steps. The optimizations are performed using the L-BFGS method [86].
In order to obtain ground states with high fidelities, following the energy minimization described above, we use the resulting wave functions as a starting point for an overlap-based optimization, where the optimizer minimizes the loss function $f=\log_{10}(1-|\bra{\Psi_{VQE}}\ket{\Psi_{DMRG}}|^{2})$ (the logarithm of the infidelity with respect to the DMRG ground state). We find that this hybrid optimization strategy yields the overall best results in terms of minimizing the energy and also producing a high-fidelity ground state. We perform ten independent optimizations following the above hybrid energy-/overlap-based minimization, and we take the state with the minimal energy among these ten as the ground state, which prevents the system from becoming stuck in local minima.
An important factor in the ground state optimization is the electron filling of the bands of the different materials studied here. $\text{Ca}_{2}\text{CuO}_{3}$ and $\text{SrVO}_{3}$ are metallic at the DFT level of theory, and we solve for their ground state at half-filling of the first electronic band included in the model. $\text{WTe}_{2}$ is predicted to be a conventional band insulator at the Kohn-Sham level of theory, we therefore solve for the ground state of this system with an initial state where the two lower-energy bands in the active space (see Fig. 3 in main manuscript) are at full-filling, and the two upper bands are empty. Since here we work in the basis of Wannier functions, we populate the Wannier states with the greatest contribution from the two lower-energy Kohn-Sham states, averaged across the Brillouin zone, as can be deduced from the Wannier rotation matrices obtained within Wannier90 [49].
In Tables 3-6 we give the VQE energies and fidelities obtained for $\text{WTe}_{2}$ and $\text{SrVO}_{3}$, ranked from best to worst from right to left.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$115.097$ eV $117.327$ eV $118.903$ eV $122.972$ eV $129.077$ eV $130.034$ eV $131.099$ eV $133.006$ eV $134.170$ eV $134.216$ eV
DMRG Energy: $115.029$ eV
Table 3: VQE energies of $\text{WTe}_{2}$.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$96.2\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Table 4: VQE fidelities of $\text{WTe}_{2}$ with respect to DMRG reference.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$-105.365$ eV $-105.363$ eV $-105.298$ eV $-105.260$ eV $-105.167$ eV $-105.120$ eV $-101.833$ eV $-84.214$ eV $-84.094$ eV $-82.473$ eV
DMRG Energy: $-105.383$ eV
Table 5: VQE energies of $\text{SrVO}_{3}$.
VQE 1 VQE 2 VQE 3 VQE 4 VQE 5 VQE 6 VQE 7 VQE 8 VQE 9 VQE 10
$31.8\%$ $11.7\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Table 6: VQE fidelities of $\text{SrVO}_{3}$ with respect to DMRG reference.
Appendix B Details of DFT and constrained RPA calculations
We perform all DFT calculations within the Quantum Espresso software package [48], within the generalized gradient approximation (GGA) of Perdew, Burke and Ernzerhof (PBE) [68]. We utilize scalar-relativistic optimized norm-conserving Vanderbilt pseudopotentials (ONCV) [87] with standard accuracy, as these are given in Pseudo Dojo [88].
For our DFT calculations on $\text{Ca}_{2}\text{CuO}_{3}$ we employ a wave function cutoff of $80$ Ry, and a $6\times 6\times 6$ k-grid. We compute the dielectric function and the Coulomb integrals of the system within RESPACK [23], using a polarizability cutoff of $7$ Ry and $100$ bands, excluding a single band crossing the Fermi level within cRPA. For $\text{WTe}_{2}$ we use a wave function cutoff of $80$ Ry, and a $6\times 6\times 1$ k-grid. We compute the dielectric function within cRPA by excluding the four bands around the Fermi level, with a polarizability cutoff of $5$ Ry and $600$ bands. For $\text{SrVO}_{3}$ we use a wave function cutoff of $81$ Ry, and a $6\times 6\times 6$ k-grid. We compute the dielectric function within cRPA by excluding the three bands crossing the Fermi level, with a polarizability cutoff of $5$ Ry and $600$ bands, yielding Coulomb parameters in close agreement to those reported previously [23].
Appendix C Downfolded Hamiltonian parameters
Here we give the parameters resulting from downfolding the electronic structure of the different materials onto the extended Hubbard Hamiltonian of eq. 1 of the main manuscript. Here we give the nearest-neighbor terms along the crystallographic direction where the maximal coupling and interactions occur, for each system.
C.1 $\text{Ca}_{2}\text{CuO}_{3}$
The hopping and Coulomb terms of this system are dominant along the crystallographic direction which aligns with chains of Cu atoms. We therefore construct a one-dimensional, single-band Hubbard model with the following parameters resulting from the Wannierization and downfolding procedures: hopping integral of $t=-0.491$ eV, on-site Coulomb interaction of $U=3.578$ eV and off-site Coulomb repulsion of $V=0.903$ eV.
C.2 $\text{WTe}_{2}$
Within the subspace of four bands around the Fermi level of $\text{WTe}_{2}$, we find the hopping term (all values in eV)
$t_{i\mathbf{R}j\mathbf{R}^{\prime}}=\begin{pmatrix}-0.201&0.178&-0.398&-0.128\\ 0.108&-0.144&0.072&-0.071\\ 0.398&0.003&0.387&0.025\\ 0.019&0.071&0.057&0.124\\ \end{pmatrix}.$ (12)
with $i=j$ the intra-band terms, and $i\neq j$ the inter-band contributions, for $\mathbf{R},\mathbf{R}^{\prime}$ corresponding to nearest neighbors.
Similarly, for the on-site Coulomb interaction
$U_{ij}=\begin{pmatrix}1.107&0.822&0.922&0.765\\ 0.822&1.095&0.760&0.684\\ 0.922&0.760&1.096&0.853\\ 0.765&0.684&0.853&1.174\\ \end{pmatrix},$ (13)
and the nearest-neighbor off-site terms:
$V_{ij}=\begin{pmatrix}0.924&0.822&0.841&0.765\\ 0.754&0.917&0.715&0.672\\ 0.841&0.760&0.855&0.853\\ 0.721&0.672&0.762&0.860\\ \end{pmatrix}.$ (14)
C.3 $\text{SrVO}_{3}$
$\text{SrVO}_{3}$ has cubic symmetry, making the Hamiltonian parameters identical along the three crystallographic axes. We find within the subspace of the three electronic bands crossing the Fermi level that we have the following intra- and inter-band terms, where all values are given in eV. For the hopping term
$t_{i\mathbf{R}j\mathbf{R}^{\prime}}=\begin{pmatrix}-0.263&0&0\\ 0&-0.263&0\\ 0&0&-0.027\\ \end{pmatrix}.$ (15)
The on-site Coulomb interaction
$U_{ij}=\begin{pmatrix}3.527&2.349&2.349\\ 2.349&3.527&2.349\\ 2.349&2.349&3.527\\ \end{pmatrix},$ (16)
and the nearest-neighbor off-site terms:
$V_{ij}=\begin{pmatrix}0.649&0.635&0.555\\ 0.635&0.649&0.555\\ 0.555&0.555&0.492\\ \end{pmatrix},$ (17)
where here too all values are given in eV.